### v1

# K–P Quantum Neural Networks

Abstract
We present an extension of K–P time-optimal quantum control solutions using global Cartan $KAK$ decompositions for geodesic-based solutions. Extending recent time-optimal constant–$\theta$ control results, we integrate Cartan methods into equivariant quantum neural network (EQNN) for quantum control tasks. We show that a finite-depth limited EQNN ansatz equipped with Cartan layers can replicate the constant–$\theta$ sub-Riemannian geodesics for K–P problems. We demonstrate how for certain classes of control problem on Riemannian symmetric spaces, gradient-based training using an appropriate cost function converges to certain global time-optimal solutions when satisfying simple regularity conditions. This generalises prior geometric control theory methods and clarifies how optimal geodesic estimation can be performed in quantum machine learning contexts.
Keywords:
Quantum control K–P problem Equivariant QNN Cartan decomposition Optimal geodesics Sub-Riemannian geometry Machine learning.
1 Introduction
Time-optimal control of quantum systems is central to many areas of quantum technology, ranging from fast gate synthesis in quantum computing to high-fidelity pulse shaping in nuclear magnetic resonance [7, 12, 5]. The K–P problem [4] is a canonical formulation of such time-optimal tasks. K–P problems involve a semisimple Lie algebra $\mathfrak{g}$ into $\mathfrak{k}\oplus\mathfrak{p}$ under a Cartan (or involution-based) decomposition [8]. The physically available (horizontal) controls come from $\mathfrak{p}$, while the compact part $\mathfrak{k}=[\mathfrak{p},\mathfrak{p}]$ must be generated indirectly via commutators. Previous work has shown that sub-Riemannian geometry on $\mathfrak{g}$ yields geodesics for locally time-optimal motion [6, 2]. Recently, a new method of optimal control was demonstrated [14] employing global Cartan $KAK$ decompositions and enforcing a constant–$\theta$ condition, one obtains an analytically solvable geodesic for quantum control problems on certain classes of Riemannian symmetric space. Other recent work has examined relaxing the requirement of full equivariance for qubit systems using variational quantum algorithms generated by horizontal elements $\mathfrak{p}$ of $KAK$ structures [16]. This partial respecting of symmetry translations is equivalent to classes of sub-Riemannian control problems invariant under translations $\mathfrak{p}$ but not generators of rotations $\mathfrak{k}$. We extend these results in two directions:
  1. 1.
    K–P QNN (EQNN) Integration.  We show how to parameterize the same Cartan-based geodesics in a neural-network ansatz that is equivariant to the underlying symmetry group. Extending existing work, we show how as with full EQNNs, composing layer-wise networks using our global method which respects the $\mathfrak{k}/\mathfrak{p}$ partition enables the networks (and importantly, their outputs) to respect sub-Riemannian symmetries. We subsequently explore how K–P respecting networks converge to approximate sub-Riemannian geodesics, without requiring the user to solve the geodesic equations symbolically.
  2. 2.
    Time-optimal QNNs  We demonstrate how QNNs integrating the K–P structure give rise to time-optimal solutions. Specifically we show (i) the existence of a solution such that finite-depth EQNNs with appropriate Cartan layers can exactly represent the constant–$\theta$ geodesics found in [14]; and (ii) the uniqueness of a solution such that that any local optimum of a suitably chosen cost function (fidelity plus sub-Riemannian penalty) may, under certain circumstances, converge to the global optimum (where target unitaries $U_{target}$ are not in the centralizer of $G$).
Section 2 reviews the K–P setup and the essential Cartan machinery for the constant–$\theta$ approach. Section 3 introduces K–P quantum neural networks sub-Riemannian layers. Section 4 states and proves the main theorems on existence and uniqueness for global optimality specific choices of target not in the centralizer of $\mathfrak{g}$. Section 5 provides examples and numerical illustrations. Finally, Section 6 offers concluding remarks on open problems.
2 Background
2.1 Cartan decompositions
Let $G$ be a connected semisimple Lie group (compact for simplicity of exposition) with Lie algebra $\mathfrak{g}$. A Cartan involution $\chi$ partitions $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$, where $\mathfrak{k}$ is the $+1$ eigenspace ($\chi(X)=+X$ for $X\in\mathfrak{k}$) and $\mathfrak{p}$ the $-1$ eigenspace ($\chi(X)=-X$ for $X\in\mathfrak{p}$). In many quantum applications, $G\subseteq U(n)$ and $\mathfrak{k}$ is the maximal compact part while $\mathfrak{p}$ is noncompact. This gives rise to Cartan commutation relations [8]:
$[\mathfrak{k},\mathfrak{k}]\subseteq\mathfrak{k},\quad[\mathfrak{p},\mathfrak{p}]\subseteq\mathfrak{k},\quad[\mathfrak{k},\mathfrak{p}]\subseteq\mathfrak{p}.$ (1)
$K=\exp(\mathfrak{k})$ is a subgroup of $G$. If $G$ is compact and semisimple then $K$ is typically a maximal torus. A typical quantum control scenario involves Hamiltonians comprising generators in $\mathfrak{p}$, while the evolution generated by the $\mathfrak{k}$ part arises via the commutators in (1). We want to implement a target $U_{\mathrm{target}}\in G$ in minimal time subject to an energy cutoff $\|H(t)\|\leq\Omega$, with $H(t)\in\mathfrak{p}$. In sub-Riemannian geometric terms, $\mathfrak{p}$ defines the horizontal distribution. $\mathfrak{k}$ is the vertical direction that can be reached by curvature forms $[\mathfrak{p},\mathfrak{p}]\subset\mathfrak{k}$ [10]. The corresponding group $KAK$ decomposition is given by $G=K\exp(\mathfrak{a})K$, where $\mathfrak{a}\subset\mathfrak{p}$ is an abelian subalgebra (maximally noncompact Cartan). Elements in $\mathfrak{a}$ typically look like $i\Theta$, with $\Theta$ real diagonal in a suitable representation. Then any $U\in G$ can be written:
$U\;=\;k\,\exp(i\,\Theta)\,c\quad\text{for some}\;k,c\in K,\;\Theta\in\mathfrak{a}.$
[14] showed that the time-optimal solutions $\gamma(t)$ can be simplified if we impose $d\Theta(t)=0$ along the path, the so-called constant–$\theta$ condition. Under mild conditions, this yields a closed-form geodesic and the minimal time $T$ is related to $|\sin(\mathrm{ad}_{\Theta})(\Phi)|$, with $\Phi\in\mathfrak{k}$ in the commutant of $\Theta$. The result in [14] can be expressed as follows.
Theorem 2.1 (Constant–$\theta$ K–P Geodesics)
Let $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ be a Cartan decomposition of a compact semisimple Lie algebra and let $\Theta\in\mathfrak{a}\cap\mathfrak{p}$ be in the noncompact Cartan subalgebra. Suppose $\Phi\in\mathfrak{k}$ commutes with $\Theta$. Then if $H(t)\in\mathfrak{p}$ saturates $\|H\|=\Omega$ and satisfies the minimal connection plus $\dot{\Theta}(t)=0$, the minimum time $T$ optimal path from $U(0)=\exp(\mathrm{i}\Theta)$ to $U(T)=\exp(-\mathrm{i}X)$ ($X\in\mathfrak{k}$ a target) has length
$\Omega\,T\;=\;\bigl{|}\sin(\mathrm{ad}_{\Theta})(\Phi)\bigr{|}$ (2)
subject to $X=(1-\cos(\mathrm{ad}_{\Theta}))(\Phi)$.
This solution leads to:
$U(t)\;=\;\exp\bigl{(}-\mathrm{i}\,\Lambda\,t\bigr{)}\,\sin(\mathrm{ad}_{\Theta})(\Phi)\,\exp\bigl{(}+\mathrm{i}\,\Lambda\,t\bigr{)}\quad\text{with}\quad\Lambda=\frac{\cos(\mathrm{ad}_{\Theta})(\Phi)}{T}.$ (3)
See [14, 13] for proofs and exposition. Theorem 2.1 shows that for certain classes of quantum control problem corresponding to symmetric spaces, there exist globally optimal controls can be solved analytically. Below, we adapt and integrate Theorem 2.1 into a quantum neural network setting.
3 K–P QNNs
EQNNs architect neural networks to respect group symmetries in ways that facilitate task optimisation [11, 15] and have shown success in quantum optimisation tasks. EQNNs have layers designed so that transformations by a group $G$ act consistently on inputs and outputs [9]. The K–P problem has an explicit decomposition $\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p}$ associated with the involution $\chi$. A natural route to building an QNN respecting K–P structure is to make the layer transformations equivariant with respect to symmetry subgroup $K$ (the subgroup generated by $\mathfrak{k}$). In this formulation, unitary conjugation $k\in K$ transforms the QNN parameters consistent with Eqn. (1) above. In the simplest sense, an K–P layer can be written:
$\mathcal{L}_{\boldsymbol{\theta}}(\rho)\;=\;\exp\bigl{(}-\mathrm{i}\,H_{\mathrm{p}}(\boldsymbol{\theta})\bigr{)}\,\rho\,\exp\bigl{(}+\mathrm{i}\,H_{\mathrm{p}}(\boldsymbol{\theta})\bigr{)},$ (4)
where $H_{\mathrm{p}}(\boldsymbol{\theta})\in\mathfrak{p}$ is restricted to the horizontal subalgebra $\mathfrak{p}$. Conjugation by $\mathfrak{k}$ then implements a vertical shift, but is consistent with the underlying symmetry (cycling generators within $\mathfrak{k}$ and $\mathfrak{p}$ respectively). The key relation is set by $[\mathfrak{p},\mathfrak{p}]\subset\mathfrak{k}$ which allows generalised rotations to be synthesised in a controlled manner using generators in $\mathfrak{p}$, we can keep the distribution structure.In K–P tasks we typically want controls only in $\mathfrak{p}$. To replicate the constant–$\theta$ geodesics of [14], we construct layer Eq. (4) in a way that fixes $\Theta$ and so that the net effect on $\mathfrak{k}$ arises from commutators $[\mathfrak{p},\mathfrak{p}]$. In [14], the minimal connection is given by:
$k^{\,\text{-}1}dk\;=\;-\cos(\mathrm{ad}_{\Theta})(dc\,c^{\,\text{-}1}),$
and hence $U(t)$ could be generated purely by $\mathfrak{p}$. We encode such constraints into the QNN layer as follows.
  1. 1.
    Initialize $\Theta$ (frozen). First, we choose an element $\Theta\in\mathfrak{a}\cap\mathfrak{p}$ (so it is in the noncompact Cartan subalgebra). In an N-qubit representation, this might be a block diagonal or simple diagonal with real entries.
  2. 2.
    Parameterize commutant $\Phi\in\mathrm{Comm}(\Theta)\cap\mathfrak{k}$. Because $\Phi$ commutes with $\Theta$, $\mathrm{Ad}_{\exp(\mathrm{i}\Theta)}(\Phi)=\Phi$.
  3. 3.
    Generate horizontal pulses. The net effect of turning on $\Phi$ plus $\sin(\mathrm{ad}_{\Theta})(\Phi)\in\mathfrak{p}$ replicates the constant–$\theta$ geodesic. The layer exponentiates $\sin(\mathrm{ad}_{\Theta})(\Phi)$ for a certain amplitude $\alpha$, while also exponentiating $\Phi$ for a turning rate $\lambda$.
  4. 4.
    Repeat in a multi-layer QNN. Several such layers can be stacked or interleaved with standard universal gates. If the objective is to achieve the final $U_{\mathrm{target}}$ with minimal $\|H\|\,T$, the network can be trained via a cost function $C=1-\mathrm{fidelity}+\kappa\cdot(\text{time penalty})$, just as in typical VQA approaches [3].
Because the layer is built from $\{\Theta,\Phi\}$ with $\Theta$ fixed, the QNN is automatically equivariant under transformations in $K$ that fix $\Theta$ (or map it to an isomorphic subalgebra). In practice, the numerical training need not solve the entire geodesic system explicitly, but the final result (provided local minima are avoided) matches the sub-Riemannian solution.
4 Existence and uniqueness of K–P circuits
We now set out results showing (1) the existence of a finite-depth EQNN with Cartan layers can represent the constant–$\theta$ K–P solution, and (2) the uniqueness of cost function minima via convergence with global minima once sub-Riemannian constraints are imposed. First, we show the existence of a finite-depth EQNN circuit for the constant-$\theta$ solution.
Theorem 4.1 (K–P QNN Circuit (Existence))
Consider a quantum system whose algebra $\mathfrak{g}$ has a Cartan decomposition $\mathfrak{k}\oplus\mathfrak{p}$. Let $\Theta\in\mathfrak{a}\cap\mathfrak{p}$, and let $\Phi\in\mathfrak{k}$ commute with $\Theta$. Then there is a finite-depth EQNN ansatz,
$U(\boldsymbol{\alpha})\;=\;\prod_{j=1}^{m}\exp\bigl{(}\mathrm{i}\,\alpha_{j}^{(\mathfrak{p})}\,Y_{j}^{(\mathfrak{p})}\bigr{)}\;\exp\bigl{(}\mathrm{i}\,\alpha_{j}^{(\mathfrak{k})}\,Y_{j}^{(\mathfrak{k})}\bigr{)},$ (5)
with $Y_{j}^{(\mathfrak{p})}\in\mathfrak{p}$, $Y_{j}^{(\mathfrak{k})}\in\mathfrak{k}$ and controls $\alpha\in\mathbb{R}^{n}$ (for $n=\dim\mathfrak{g}$), that can exactly realize the constant–$\theta$ solution in Theorem 2.1 for any choice of $\Theta,\Phi$. In particular, $U(\boldsymbol{\alpha}^{*})=U(T)$ from Theorem 2.1 for some $\boldsymbol{\alpha}^{*}$.
Proof
As $\Phi\in\mathrm{Comm}(\Theta)$, then $\mathrm{Ad}_{\exp(\mathrm{i}\Theta}(\Phi)=\Phi$, and $\mathrm{ad}_{\Theta}(\Phi)=[\Theta,\Phi]$ remains in $\mathfrak{p}$. The Trotter expansions of $\exp(\sin(\mathrm{ad}_{\Theta})(\Phi))$ can be compiled into a finite product of exponentials in $\mathfrak{p}$ and $\mathfrak{k}$. Letting $Y^{\mathfrak{p}}_{1}=[\Theta,\Phi]\in\mathfrak{p}$, we see that $\sin(\mathrm{ad}_{\Theta})(\Phi)$ is a linear combination of repeated commutators of $Y^{\mathfrak{p}}_{1}$ with $\Theta$. Repeated commutators remain in $\mathfrak{p}$ or $\mathfrak{k}$ by the standard Cartan relations. Hence, we can approximate $\exp\bigl{(}\mathrm{i}\,\sin(\mathrm{ad}_{\Theta})(\Phi)\bigr{)}$ as a product of exponentials of $\mathfrak{p}$ or $\mathfrak{k}$. Appending a single factor $\exp(\mathrm{i}\,\Lambda\,t)$ in $\mathfrak{k}$ recovers the $\cos(\mathrm{ad}_{\Theta})(\Phi)$ term. Thus, (5) suffices to represent $U(T)$. The condition that $\boldsymbol{\alpha}\mapsto(Y_{j}^{(\mathfrak{p})},Y_{j}^{(\mathfrak{k})})$ behaves in an equivariant manner (i.e. transforms consistently under $k\in K$) follows from choosing the $Y_{j}$ in each layer to be sums of $\mathfrak{p}$ or $\mathfrak{k}$ sub-blocks, respecting $\chi$. ∎
The control parameters $\alpha$ may be learnt and optimised according to a typical VQA optimisatoin algorithm. Next, we show that any local optimum of a standard cost function in the EQNN approach must coincide with the global sub-Riemannian geodesic solution.
Theorem 4.2 (K–P Stationary Points: Uniqueness)
Let  $U(\boldsymbol{\alpha})$ be a finite-depth circuit of the form (5), where each layer is generated by operators in $\mathfrak{p}$ (the “horizontal” subalgebra) with or without additional commutator-generated rotations from $\mathfrak{k}$. Define the cost function
[…]
$\nabla_{\!\boldsymbol{\alpha}}\,C(\boldsymbol{\alpha}^{*})\;=\;0\quad\text{and}\quad\|H_{\mathfrak{p}}(t)\|\leq\Omega\;\,\forall\,t\,.$ (7)
Then $U(\boldsymbol{\alpha}^{*})$ coincides with the constant–$\theta$ geodesic from Theorem 2.1 and is (globally) time-optimal under the K–P constraints.
Proof
[…]
Theorems 4.1 and 4.2 show that (for certain choices of subgroup $K$ (and thus symmetric space) that K–P QNNs can be made both complete and globally convergent: (i) the finite-depth circuit is guaranteed to express a minimal-time path; and (ii) gradient-based optimisation (with an appropriate cost) can avoid spurious local minima. We briefly illustrate how K–P QNNs with a can discover the same time-optimal solution as an analytic or geometric approach.
5 Evaluation: $\Lambda$-systems and $SU(3)$
Consider the three-level $\Lambda$-system studied in [14] where results from [1] were reproduced using the constant-$\theta$ method. $\mathrm{SU}(3)$, has a representation in terms of Gell-Mann generators which can be decomposed into a horizontal subalgebra $\mathfrak{p}\subset\mathfrak{g}$ (those used as direct controls) and a vertical subalgebra $\mathfrak{k}\subset\mathfrak{g}$ (generated indirectly via commutators). The $\Lambda$ system is a three-level model with horizontal transitions $\mathfrak{p}$ coupling two ground states to an excited state. To evaluate our K–P QNN, we construct python code to test its efficacy at reproducing the $\Lambda$-system optimisation results in [14]. We select:
$\mathfrak{k}\;=\;\mathrm{span}\bigl{\{}\,{-i\lambda_{3}},\,{-i\lambda_{6}},\,{-i\lambda_{7}},\,{-i\lambda_{8}}\bigr{\}},\quad\mathfrak{p}\;=\;\mathrm{span}\bigl{\{}\,{-i\lambda_{1}},\,{-i\lambda_{2}},\,{-i\lambda_{4}},\,{-i\lambda_{5}}\bigr{\}}.$
Arrays store these respective matrices. We construct a finite-depth product of exponentials:
$U(\boldsymbol{\alpha})\;=\;\prod_{j=1}^{L}\exp\!\Bigl{(}\mathrm{i}\!\sum_{a\in\mathfrak{k}}\alpha_{j,a}^{(\mathfrak{k})}\,Y_{a}\Bigr{)}\;\exp\!\Bigl{(}\mathrm{i}\!\sum_{b\in\mathfrak{p}}\alpha_{j,b}^{(\mathfrak{p})}\,X_{b}\Bigr{)}.$
Here, $\{Y_{a}\}\subset\mathfrak{k}$ and $\{X_{b}\}\subset\mathfrak{p}$ are the basis elements. We implement this as the function circuit_forward, which sequentially multiplies each layer’s matrix exponential (see Theorem 4.1). From Theorem 4.2 (cf. Eq. (6)), we define a cost that has two main terms:
$C(\boldsymbol{\alpha})\;=\;\bigl{[}\,1-\mathrm{Fid}\bigl{(}U(\boldsymbol{\alpha}),\,U_{\mathrm{target}}\bigr{)}\bigr{]}\;+\;\gamma\,\bigl{(}\text{path\_length}(\boldsymbol{\alpha})\bigr{)}.$
In the code, we approximate the sub-Riemannian path length by the sum of the $\ell_{2}$-norms of the horizontal parameters $\alpha^{(\mathfrak{p})}$ encoded in path\_length. This is a simplified version of $\int_{0}^{T}\|H_{\mathfrak{p}}(t)\|\mathrm{d}t$, sufficient to demonstrate the principle of penalizing the magnitude of $\mathfrak{p}$-controls. We apply a straightforward gradient descent (using JAX auto-differentiation). By Theorem 4.2, no spurious local minima exist if $\boldsymbol{\alpha}$ saturates the bracket generation assumptions and $\|H(t)\|\leq\Omega$. Convergence to fidelity $\approx 1$ and small path length indicates that the learned solution reproduces the sub-Riemannian geodesic described in the main text.
Figure 1: K–P QNN Loss function (Eqn. (6) showing convergence between $U(\alpha)$ and $U_{target}$. Convergence shows the K–P QNN learning the optimal unitary parametrised by controls $\alpha$ (reproducing targets in [1] and [14]).
As shown in Fig. (1), after sufficient epochs, the cost function converges close to zero, and the final unitary $U(\boldsymbol{\alpha}^{*})$ attains $\mathrm{Fid}\approx 1$ (specifically 0.9998576). Hence, the constant–$\theta$ time-optimal solution $\exp\bigl{(}-\tfrac{\mathrm{i}\pi}{4}\,\lambda_{6}\bigr{)}$ is accurately reconstructed, in agreement with the analytical results (cf. Section 4 and Eq. (2)). Thus the K–P QNN approach recovers the same geodesic solution from a purely data-driven perspective. The repository is available at https://github.com/eperrier/k-p_qnn.
6 Conclusions and Outlook
We have shown how K–P QNNs, a form of EQNNs can be naturally constructed using the constant–$\theta$ Cartan decomposition approach to the K–P problem. The synergy arises from the geometric consistency: EQNN layers that keep $\Theta$ fixed and use pulses in $\mathfrak{p}$ effectively trace out the same sub-Riemannian geodesic described by [14]. Our results show that local optimality in a typical QNN variational cost function indeed implies global time-optimality for certain classes of quantum control problems on symmetric spaces $G/K$. Limitations of our method include those set out in [14], particularly that time-optimal sequences are found for only certain targets in $\mathfrak{g}$. Future research directions building on this work may consider:
  • incorporating noise and decoherence by letting the cost function measure fidelity under realistic noise channels;
  • extending to non-compact or indefinite metrics (e.g. involving indefinite Killing forms or non-compact groups); or
  • experimental demonstration via implementation of EQNNs using NISQ superconducting devices.
Our work contributes to the growing literature connecting Cartan-based geodesic solutions to quantum machine learning protocols.
{credits}
6.0.1 Acknowledgements
The author thanks Chris Jackson for his considerable mentorship and discussions on the topic. This work was independently funded by the author.
[…]
The author declares that they have no conflicts of interest.